An aid in the mathematical department office gets paid $3 per hour for clerical work and $8 per hour for tutoring. If she wants to work a total of 20 hours and earn at least $145, what is the maximum number of hours X she can spend on clerical work?
Answer by stanbon(75887)

An aid in the mathematical department office gets paid $3 per hour for clerical work and $8 per hour for tutoring. If she wants to work a total of 20 hours and earn at least $145, what is the maximum number of hours X she can spend on clerical work?
------------------
Let number of hours of clerical work be "x" ; pay for this is 3x dollars
Number of hours of tutoring is "20-x"; pay for this is 8(20-x)= 160-8x dollars
-----------------------
EQUATION:
pay + pay =145 dollars
3x + 160-8x = 145
-5x = -15
x = 3 (number of clerical hours)
20-x = 17 (number of tutoring hours)
================
Cheers,
Stan H.